# MTIL: Encoding Full History with Mamba for Temporal Imitation Learning

### v1

@@ 4.2 LIBERO Benchmark
Table 2: Lifelong Learning Performance on LIBERO (EWC Strategy).
@@
FWT($\uparrow$): Forward Transfer, NBT($\downarrow$): Backward Transfer, AUC($\uparrow$): Area Under Curve. EWC strategy results averaged over 3 seeds (100, 200, 300) at 50 epochs. Baselines from [51]. Short-history (10-step,similar performance for 20/50 steps) and full-history result shown.
4.3 Robomimic (Low-Dimensional State)
To assess MTIL’s ability to model complex dynamics from proprioceptive states alone, we evaluated it on low-dimensional Robomimic tasks [12], comparing against prior methods including Diffusion Policy [7], Even without visual input, MTIL leveraging full history consistently surpasses previous state-of-the-art methods across the evaluated tasks.(Table 3).
Table 3: Success Rates on Robomimic Low-Dimensional Tasks.
@@ 4.3 Robomimic (Low-Dimensional State)
Method ph mh ph mh ph mh ph mh ph mh
LSTM-GMM [12] 1.00 0.93 1.00 0.81 0.95 0.59 0.76 0.20 0.67 0.31
IBC [22] 0.79 0.15 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00
BET [26] 1.00 0.99 1.00 0.90 0.76 0.43 0.38 0.06 0.58 0.20
DiffusionPolicy-C [7] 1.00 0.97 1.00 0.96 0.97 0.82 0.94 0.46 0.50 0.30
DiffusionPolicy-T [7] 1.00 1.00 1.00 0.94 1.00 0.81 1.00 0.35 1.00 0.87
MTIL (10-step) 1.00 0.94 1.00 0.83 0.96 0.61 0.78 0.22 0.70 0.33
MTIL (Full History) 1.00 1.00 1.00 0.96 0.98 0.83 1.00 0.48 1.00 0.89
An MTIL variant using only a 10-step history performs comparably, albeit slightly better, than LSTM-GMM [12], indicating limited gains from short recurrent contexts here. The full-history configuration, however, demonstrates the strength of Mamba’s architecture within MTIL for capturing intricate state transition dynamics and temporal dependencies inherent in manipulation, proving effective beyond visuomotor control.
4.4 Real-World Dual-Arm Tasks
@@
To validate MTIL in complex physical environments, we designed challenging tasks on a dual UR3 platform equipped with custom 2-finger grippers and four cameras providing multi-view observations (Figure 3). We compare MTIL (using DINOv2 backbone and full history) against ACT trained on identical demonstration data (100 demos per task).
@@
We designed this task (visualized in Figure 4) specifically to challenge Markovian policies by requiring long-term memory, a scenario where methods like ACT often fail. The four stages involve: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Right arm inserts Tube1 into Tube2, (4) Right arm inserts Tube1 into Tube3. Critically, executing Stage 3 correctly necessitates recalling the completion of previous stages, as intermediate observations can be ambiguous. Table 4 details the stage-wise success rates. MTIL, leveraging its full history state, successfully completes the entire sequence with high probability. In stark contrast, ACT, reliant on immediate context, is confounded by the temporal ambiguity; it frequently attempts Stage 4 directly after Stage 2, failing to execute the required sequence correctly and resulting in zero success for completing Stage 3, Stage 4, and the overall task. This outcome underscores the limitations of short-history approaches and validates the imperative of encoding complete history for reliably executing temporally complex manipulation sequences.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
ACT 80 64 0 0 0
MTIL 94 80 62 54 54
Figure 4: Figure & Table 4: Sequential Insertion task stages (left) and success rates (%, right). MTIL successfully completes the sequence while ACT fails due to temporal ambiguity between stages 3 and 4.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
ACT 80 64 50 32 32
MTIL 94 80 74 62 62
Figure 5: Figure & Table 5: Coordinated Pouring task stages (left) and success rates (%, right). MTIL achieves higher success and smoother execution compared to ACT.
@@
This task (Figure 5) assesses precise bimanual coordination over a longer sequence: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Left arm grasps Tube2, (4) Right arm pours water from Tube1 into Tube2. While less susceptible to the specific ambiguity of the insertion task, it still requires accurate, temporally coordinated actions. Table 5 shows that although both methods achieve non-zero success, MTIL consistently outperforms ACT across the stages, resulting in a higher overall success rate and exhibiting notably smoother execution trajectories.
5 Conclusion
@@
The trajectory of intelligence is intrinsically linked to the capacity for memory – the ability to weave the tapestry of past experiences into the fabric of present action. This work confronts a central limitation in contemporary imitation learning: the prevalent reliance on the Markovian assumption, which often reduces complex sequential behaviors to mere reactions to the immediate sensory world, neglecting the crucial dimension of history. We introduced Mamba Temporal Imitation Learning (MTIL), a paradigm shift that embraces the power of memory by leveraging the recurrent state dynamics inherent within the Mamba architecture. MTIL demonstrates that by encoding the full history of observations into a compressed, evolving state representation, robots can transcend the confines of the present moment. This comprehensive temporal context allows MTIL to effectively disambiguate perception and unlock the execution of intricate, state-dependent sequential tasks previously challenging for established methods like ACT and Diffusion Policy, as validated through extensive experiments across diverse simulated and real-world benchmarks. Our findings not only showcase the significant performance and efficiency gains afforded by MTIL but, more profoundly, underscore the essential role of history in bridging the gap between perception and intelligent, temporally coherent action in robotics. By demonstrating the efficacy of State Space Models in capturing the long flow of time, this work illuminates a promising pathway towards building robotic agents capable of deeper understanding and more sophisticated interaction with the world.
Limitations
While MTIL demonstrates significant advantages, we acknowledge certain limitations. Firstly, our current training strategy relies on sequential processing via Mamba’s ‘step‘ function(as detailed in Section 3.3). This is necessitated by the significant GPU memory demands of processing entire long sequences of high-dimensional observations concurrently, a challenge for current hardware. Unlike methods operating under the Markov assumption which might permit data shuffling, MTIL must preserve temporal order, inherently limiting parallelization compared to approaches like windowed Transformers and resulting in slower wall-clock training times per epoch—a trade-off for enabling full history modeling. Exploring techniques like gradient checkpointing or specialized hardware optimizations presents an avenue to mitigate this training speed limitation. Secondly, a potential failure mode, common to many imitation learning approaches, stems from distributional shift between training demonstrations and deployment scenarios. MTIL’s generalization capability to substantially novel environments or configurations may be limited, potentially leading to performance degradation when facing significant train-test mismatch.Lastly, while Mamba’s state aims to compress history effectively, its fidelity over ultra-long sequences (e.g., tens of thousands of steps) remains to be rigorously verified. Furthermore, performance can exhibit sensitivity to hyperparameter choices within the Mamba architecture and training regime, necessitating careful tuning. Addressing these aspects, alongside further theoretical investigation into the state representation limits, remains important future work.
@@ Acknowledgments
This work is supported by the Joint Funds of the National Natural Science Foundation of China (Grant No. U [^2] A20208), the Natural Science Foundation Innovation Group Project of Hubei Province (Grant No. 2022CFA018), and the Key Research and Development Program of Guangdong Province (Grant No. 2022B0202010001-2).
Appendix A Appendix
A.1 ACT Benchmark Experimental Setup
Tasks and Data.
We evaluated MTIL on the two simulated tasks introduced in the ACT paper [5]: ”Cube Transfer” and ”Bimanual Insertion”. For both tasks, we used a dataset of 100 expert demonstration trajectories, automatically generated using scripted policies provided in the original ACT repository. The typical trajectory length for these tasks is around 400 steps.
MTIL Configuration.
The main MTIL model (”MTIL (Full)”) used full history encoding, processing the entire trajectory sequentially during training as described in Algorithm 1. We used an action chunk prediction size of $K=50$. The model was trained using the AdamW optimizer [53] with an initial learning rate of $2\times 10^{-4}$. A cosine learning rate decay schedule was applied over the training epochs. For the short-history ablation (”MTIL (10-step Hist.)”), the same architecture and training setup were used, but the hidden state was reset every 10 steps, limiting the effective history length.
Backbones and Feature Processing.
We experimented with two visual backbones to compare performance and isolate the effect of history modeling:
  • ResNet18 (for fair comparison with ACT): Input RGB images (480x640x3) from multiple cameras were processed by separate ResNet18 [50] instances. Following the ACT pipeline, we extracted feature maps from an intermediate layer, resulting in 15x20x512 feature maps per camera. We applied learned positional encodings to these features. To enhance feature representation, we incorporated a Convolutional Block Attention Module (CBAM) [54] after the ResNet features, applying both channel and spatial attention mechanisms. The resulting features were flattened and projected using a linear layer to a common embedding dimension.
  • DINOv2 (ViT-L/14, for best performance): Input RGB images (480x640x3) were fed into a pre-trained DINOv2 model with a ViT-L/14 architecture [49]. The weights of the DINOv2 model were kept frozen during training. We extracted features from the output of the penultimate transformer block (specifically, ‘blocks[-4]‘ in common implementations), yielding a feature map of size approximately 34x45x1024 per camera (exact dimensions depend slightly on implementation details of patch embedding and image size). These high-dimensional features were then flattened and projected via a linear layer to the common embedding dimension.
Multi-Modal Fusion.
The projected visual features from all cameras were concatenated. Low-dimensional robot state information (e.g., joint positions, gripper width/status) was also concatenated into a vector. To effectively integrate these modalities, we employed a cross-modal attention mechanism: the concatenated visual features served as query, while the low-dimensional state vector served as the keys and values. The output of this attention mechanism was then passed through a final linear layer to project it to the input dimension ($D$) required by the Mamba-2 blocks.
Mamba-2 Architecture.
The core sequence modeling was performed by a stack of Mamba-2 blocks. Key hyperparameters such as the model dimension ($d_{model}$), state expansion factor, state dimension ($d_{state}$), and the number of layers were tuned based on validation performance. Specific values are provided in the codebase and Table 6.
A.2 LIBERO Benchmark Experimental Setup
Goal and Setup.
The experiments on LIBERO [51] aimed to evaluate MTIL’s performance in a challenging lifelong learning setting, focusing on knowledge transfer and catastrophic forgetting. To ensure fair comparison with the baselines reported in [51], we adhered closely to their experimental protocol.
Backbones and Prediction.
We used the standard ResNet and ViT [55] backbones provided within the official LIBERO benchmark codebase. We did *not* use the DINOv2 backbone in these experiments to maintain comparability. Furthermore, following the LIBERO setup for RNN/Transformer baselines, MTIL was configured to predict only the action for the next single step ($K=1$), rather than action chunks.
Language Encoder and Fusion.
Task descriptions were encoded using the standard BERT [56] encoder, consistent with the LIBERO baselines. For fusion, simple concatenation of visual features, state information, and language embeddings followed by linear projection was used, matching the baseline setup. No cross-modal attention mechanism was employed here.
Lifelong Learning Strategy: EWC.
We employed the Elastic Weight Consolidation (EWC) [52] strategy to mitigate catastrophic forgetting as tasks were learned sequentially. EWC works by calculating the Fisher Information Matrix (FIM) for the parameters of the model with respect to previously learned tasks. During training on a new task, a quadratic penalty term is added to the loss function, discouraging changes to parameters deemed important for past tasks (i.e., those with high Fisher information). Rationale for EWC with MTIL: We selected EWC not only because it is a standard lifelong learning technique used in the benchmark, but also due to its theoretical synergy with MTIL’s state-based history encoding. The Mamba-2 blocks contain parameters that govern the selective state updates ($\Delta_{t},\mathbf{B}_{t},\mathbf{C}_{t}$) and state transitions ($\mathbf{A}_{t}$). These parameters are crucial for capturing the temporal dynamics and historical context learned from demonstrations. By applying EWC, we hypothesize that the FIM helps identify and protect those parameters within the Mamba architecture that are most critical for encoding the temporal patterns learned in previous tasks. This allows MTIL to retain its ability to leverage history effectively for older tasks (leading to low NBT) while still adapting its state dynamics to accommodate new tasks (enabling high FWT).
Metrics Definition.
We report the standard LIBERO lifelong learning metrics, averaged over 3 random seeds after training on all tasks in a suite for 50 epochs using the EWC strategy:
  • FWT (Forward Transfer, $\uparrow$): Measures the influence of having learned tasks $11,...,i-$ 1 on the initial performance of learning task $i$, compared to learning task $i$ from scratch. It is calculated as:
    $FWT=\frac{1}{N-1}\sum_{i=2}^{N}(A_{i-1,i}-A_{base,i})$
    where $A_{i-1,i}$ is the accuracy on task $i$ evaluated *immediately after* training on task $i-1$ (before training on task $i$), and $A_{base,i}$ is the accuracy on task $i$ when trained in isolation (from scratch). Higher values indicate better positive transfer. (Note: LIBERO paper definition might vary slightly, we follow the standard interpretation of immediate transfer).
  • NBT (Negative Backward Transfer, $\downarrow$): Measures the average drop in performance on previously learned tasks after learning subsequent tasks (i.e., forgetting). It is calculated as:
    $NBT=\frac{1}{N-1}\sum_{i=1}^{N-1}(\max(0,A_{i,i}-A_{N,i}))$
    where $A_{i,i}$ is the final accuracy on task $i$ after training on task $i$, and $A_{N,i}$ is the accuracy on task $i$ after training on all $N$ tasks. Lower values indicate less forgetting.
  • AUC (Area Under the learning Curve, $\uparrow$): Represents the overall performance across all tasks throughout the lifelong learning process. It is typically calculated as the average accuracy across all tasks, evaluated after the model has finished training on the final task $N$:
    $AUC=\frac{1}{N}\sum_{i=1}^{N}A_{N,i}$
    Higher values indicate better overall performance retention and learning.
A.3 Robomimic Benchmark Experimental Setup
Goal and Setup.
The experiments on Robomimic [12] aimed to evaluate MTIL’s ability to model complex manipulation skills using only low-dimensional proprioceptive state information (e.g., joint positions, velocities, end-effector pose, gripper status), without any visual input.
Input and Settings.
We used the standard low-dimensional state datasets provided by Robomimic for tasks like Lift, Can, Square, Transport, and Tool Hang, considering both proficient-human (ph) and multi-human (mh) data sources. To ensure fair comparison, we followed the evaluation protocol and data splits used in prior work, particularly matching the setup for the LSTM-GMM baseline [12] where applicable.
Key Difference and Action Output.
The primary difference between MTIL and the LSTM-GMM baseline lies in the history modeling. While LSTM-GMM uses LSTMs, often trained on shorter sequences (e.g., 10-50 steps) in practice [12], MTIL (Full History) processes the entire demonstration sequence using its Mamba-based recurrent state. To isolate the effect of the sequence modeling backbone, we configured MTIL to use a Gaussian Mixture Model (GMM) head for predicting the action distribution at each step, mirroring the output structure of LSTM-GMM. The number of mixture components was treated as a hyperparameter.
MTIL Architecture.
A stack of Mamba-2 blocks was used to process the sequence of low-dimensional state vectors. The specific hyperparameters (e.g., $d_{model}$, $d_{state}$, number of layers) were optimized for these tasks (See details in Table 6).
A.4 Real-World Experimental Setup
Figure 6: Details of the experimental hardware, showing the dual-arm robot setup (UR3) with the task workspace, and the specific custom gripper, wrist camera(D405), Kinect camera and RealSense D435 utilized.
Platform.
Experiments were conducted on a physical setup consisting of two Universal Robots UR3 arms mounted on a table. Each arm was equipped with a custom 3D-printed 2-finger parallel gripper actuated by a Dynamixel servo motor. The sensory input comprised multi-view images from four cameras: one Intel RealSense D435i providing a side view, one Microsoft Azure Kinect providing a top-down view, and two Intel RealSense D405 cameras mounted on the wrists of the robots for close-up views (see Figure  6).
MTIL Configuration.
For the real-world tasks (Sequential Insertion, Coordinated Pouring), we used the full MTIL configuration designed for robust performance: full history encoding via sequential training (Algorithm 1), the DINOv2 (ViT-L/14) backbone with frozen weights for visual feature extraction, and prediction of action chunks with $K=50$. Multi-modal fusion using cross-attention between visual features and low-dimensional state (joint positions, gripper state) was employed, identical to the setup described in Appendix A.1. Specific Mamba-2 architecture parameters were consistent with those used in the ACT benchmark simulation experiments with DINOv2.
Data Collection.
For each of the two tasks, we collected 100 expert demonstrations using a custom teleoperation system that use spacemouse. The system allowed a human operator to control the dual arms simultaneously via input devices, recording the robot states and camera images at each timestep.The trajectory length for the Sequential Insertion task is around 500 steps while the trajectory length for the Coordinated Pouring task is around 1000 steps.
Inference Details: Chunking and Temporal Aggregation.
During policy execution on the real robot (inference time, Algorithm 2), we applied both action chunking (predicting $K=50$ steps ahead) and temporal aggregation. Temporal aggregation was implemented using a weighted averaging scheme over overlapping predicted chunks. Specifically, the action $a_{t}^{\text{final}}$ executed at time $t$ was computed as a weighted average of predictions made for time $t$ from chunks initiated at previous times $j$ (where $j\leq t<j+K$). We used exponentially decaying weights $W=\{w_{0},...,w_{K-1}\}$ where $w_{k}\propto\gamma^{k}$ for some decay factor $\gamma<1$, giving higher importance to predictions made closer to the current time step (i.e., smaller $k=t-j$). Rationale: Applying chunking and temporal aggregation is a common practice in methods like ACT [5] and has been shown to significantly improve the smoothness and stability of executed robot motions by reducing jitter [19, 20]. We employed this technique primarily for two reasons: (1) To ensure a fair comparison with ACT, which also utilizes this strategy. (2) To achieve better practical performance and robustness during real-world execution, as smoother actions are generally preferable. While MTIL’s inherent temporal modeling capability provides a strong foundation for coherent action sequences, the temporal aggregation acts as an additional filter that further enhances stability, leading to more reliable task execution.
A.5 Hyperparameter Summary
Table 6: Key MTIL Hyperparameters Across Benchmarks.
Parameter ACT Sim LIBERO Robomimic Real World
Backbone ResNet18 / DINOv2 ResNet / ViT N/A (State) DINOv2
History Full / 10-step Full / 10-step Full / 10-step Full
Chunk Size (K) 50 1 1 50
Mamba $d_{model}$ 2048 256 64 2048
Mamba $d_{state}$ 512 64 16 512
Mamba Layers 4 1 4 4
Optimizer AdamW AdamW AdamW AdamW
Initial LR 2e-4 1e-4 1e-4 2e-4
LR Schedule Cosine Cosine Cosine Cosine
Cross-Attention Yes No N/A Yes
Output Head Linear Linear GMM Linear
Temporal Aggregation Yes No No Yes

### v3

@@ IV-B LIBERO Benchmark
TABLE II: Lifelong Learning Performance on LIBERO (EWC Strategy).
@@
FWT($\uparrow$): Forward Transfer, NBT($\downarrow$): Negative Backward Transfer (should be Backward Transfer, if it’s negative it’s good), AUC($\uparrow$): Area Under Curve. EWC strategy results averaged over 3 seeds (100, 200, 300) at 50 epochs. Baselines from [52]. Short-history (10-step,similar performance for 20/50 steps) and full-history result shown.
IV-C Robomimic (Vision-based Policy)
TABLE III: Behavior Cloning Benchmark (Visual Policy) on Robomimic. As per the original dataset, results are reported as final success rates.
@@ IV-C Robomimic (Vision-based Policy)
ph mh ph mh ph mh ph mh ph
LSTM-GMM [29] 1.00 1.00 1.00 0.98 0.82 0.64 0.88 0.44 0.68
IBC [12] 0.94 0.39 0.08 0.00 0.03 0.00 0.00 0.00 0.00
DiffusionPolicy-C 1.00 1.00 1.00 1.00 0.98 0.98 1.00 0.89 0.95
DiffusionPolicy-T 1.00 1.00 1.00 1.00 1.00 0.94 0.98 0.73 0.76
MTIL (10-step) 1.00 1.00 1.00 0.99 0.87 0.65 0.92 0.52 0.72
MTIL (Full) 1.00 1.00 1.00 1.00 1.00 0.96 1.00 0.91 0.97
To assess MTIL’s ability to handle high-dimensional visual inputs, we evaluated it on the vision-based Robomimic tasks [12]. As shown in Table III, MTIL (Full) significantly outperforms all baselines, including the strong DiffusionPolicy variants. Notably, MTIL (10-step) offers only a marginal improvement over the LSTM-GMM baseline, highlighting that a short-history SSM is insufficient. The substantial performance gain of MTIL (Full) underscores its superior capability in leveraging full spatio-temporal context from visual data. This can be attributed to its nature as a learned dynamical system; the recurrent state $h_{t}$ acts as an implicit world model, tracking not just object locations but also their latent physical states (e.g., momentum, contact stability) over time, which is crucial for complex manipulation.
IV-D Real-World Dual-Arm Tasks
@@
TABLE IV: Sequential Insertion Success Rates (%), averaged over 50 roll-outs.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
[…]
Figure 4: Sequential Insertion task stages (left panel of figure) and success rates (right panel, Table IV). MTIL successfully completes the sequence while ACT fails due to temporal ambiguity between stages 3 and 4.
TABLE V: Coordinated Pouring Success Rates (%), averaged over 50 roll-outs.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
[…]
Figure 5: Coordinated Pouring task stages (left panel of figure) and success rates (right panel, Table V). MTIL achieves higher success and smoother execution compared to ACT.
To validate MTIL in complex physical environments, we designed challenging tasks on a dual UR3 platform equipped with custom 2-finger grippers and four cameras providing multi-view observations (Figure 3). We compare MTIL (using DINOv2 backbone and full history) against ACT trained on identical demonstration data (100 demos per task). All real-world results are averaged over 50 evaluation roll-outs for the best checkpoint from each of the 3 seeds.
@@
We designed this task (visualized in Figure 4) specifically to challenge Markovian policies by requiring long-term memory, a scenario where SOTA methods like ACT often fail. The four stages involve: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Right arm inserts Tube1 into Tube2, (4) Right arm inserts Tube1 into Tube3. Critically, executing Stage 3 correctly necessitates recalling the completion of previous stages, as intermediate observations can be ambiguous. Table IV details the stage-wise success rates. MTIL, leveraging its full history state, successfully completes the entire sequence with high probability. In stark contrast, ACT, reliant on immediate context, is confounded by the temporal ambiguity, as the observations after completing Stage 2 can be identical with completing Stage 3, making it indistinguishable for policies relying solely on current or short-term history. AS a result, it frequently attempts Stage 4 directly after Stage 2, failing to execute the required sequence correctly and resulting in zero success for completing Stage 3, Stage 4, and the overall task. This outcome underscores the limitations of short-history approaches and validates the imperative of encoding complete history for reliably executing temporally complex manipulation sequences.
@@
This task (Figure 5) assesses precise bimanual coordination over a longer sequence: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Left arm grasps Tube2, (4) Right arm pours water from Tube1 into Tube2. While less susceptible to the specific ambiguity of the insertion task, it still requires accurate, temporally coordinated actions. Table V (within Figure 5) shows that although both methods achieve non-zero success, MTIL consistently outperforms ACT across the stages, resulting in a higher overall success rate and exhibiting notably smoother execution trajectories.
V Conclusion
@@
The trajectory of intelligence is intrinsically linked to the capacity for memory – the ability to weave the tapestry of past experiences into the fabric of present action. This work confronts a central limitation in contemporary imitation learning: the prevalent reliance on the Markovian assumption, which often reduces complex sequential behaviors to mere reactions to the immediate sensory world. We introduced Mamba Temporal Imitation Learning (MTIL), a new paradigm that embraces the power of memory by leveraging the recurrent state dynamics inherent within the Mamba architecture. We posit that MTIL represents a practical and powerful synthesis of concepts from World Models and Dynamical Systems. By encoding the full history of observations into a compressed, evolving state representation, MTIL learns an implicit, action-oriented world model. This comprehensive temporal context allows MTIL to effectively disambiguate perception and unlock the execution of intricate, state-dependent sequential tasks previously challenging for established methods. Our findings not only showcase the significant performance and efficiency gains afforded by MTIL but, more profoundly, underscore the essential role of history in bridging the gap between perception and intelligent action. By demonstrating the efficacy of SSMs in capturing the long flow of time in a computationally feasible manner, this work illuminates a promising pathway towards building robotic agents capable of deeper understanding and more sophisticated interaction with the world.
@@ Acknowledgments
This work was supported by the Joint Funds of the National Natural Science Foundation of China (Grant No. U [^2] A20208), the Natural Science Foundation Innovation Group Project of Hubei Province (Grant No. 2022CFA018), and the Key Research and Development Program of Guangdong Province (Grant No. 2022B0202010001-2).